OpenStudy (anonymous):

if f(x)= 3/5x and g(x)= -3x/2, find f(x)+g(x)

OpenStudy (solomonzelman):

I will give you an example. \(\large\color{black}{ f(x)=3x+2 \\g(x)=-4x+3 }\) \(\large\color{black}{ f(x)+g(x)=(3x+2)+(-4x+3) }\) \(\large\color{black}{ f(x)+g(x)=3x+2-4x+3 }\) \(\large\color{black}{ f(x)+g(x)=-x+5 }\)

OpenStudy (solomonzelman):

so you are just adding what f(x) is equal to, and what f(x) is equal to, together. need more help, ask.